OpenStudy (bibby):

Am I setting this up correctly? A/B are using the Diffie-Hellman Key Exchange Protocol with public modulus N = 1601 and public base e = 23. A sends B 1596 and B sends A 403.

OpenStudy (bibby):

\(23^x \mod 1601 \equiv 1596\\23^x \mod1501 \equiv 403\)

OpenStudy (rational):

you want to figure out the secret key ?

OpenStudy (bibby):

yeah, I'm getting different numbers when I try setting it up this way

OpenStudy (bibby):

using this as a reference

OpenStudy (bibby):

oh nevermind, I got it. 1028

OpenStudy (rational):

how ?

OpenStudy (bibby):

I kinda cheated using wolfram. \(23^x \mod 1601 \equiv 403 \implies x=1600n+191\) 1596^191 mod 1601 = 1028 if you want an algorithmic approach, we're meant to be using pohlig-hellman or baby sttep giant step

OpenStudy (bibby):

it goes the same way in reverse. \(23^x \mod 1601 \equiv 1596 \longrightarrow x=1600n+76\) 403^76 mod 1601 = 1028

OpenStudy (rational):

1596^x = 403^y mod 1601 right ?

OpenStudy (bibby):

correct

OpenStudy (rational):

yes so the secret key is 1028

OpenStudy (bibby):

Thanks rational :D